OpenStudy (vishweshshrimali5):

Find the largest k such that \(\large{3^{11}}\) is the sum of k consecutive positive integers.

OpenStudy (vishweshshrimali5):

Okay, I tried in the following method: Let the integers be n, n+1, n+2, ..., n+k-1. Then, \[\large{\text{Sum of k consecutive positive integers} = \cfrac{(n+k-1)(n+k) + (n-1)n}{2}}\]

OpenStudy (vishweshshrimali5):

@ganeshie8

OpenStudy (vishweshshrimali5):

\[\large{\text{Sum of k consecutive positive integers} = \cfrac{(n+k-1)(n+k)}{2} +}\] \(\large{\cfrac{(n-1)n}{2}}\)

OpenStudy (vishweshshrimali5):

Now I tried to equate it to \(\large{3^{11}}\). But, where I am stuck is how I am going to calculate 2 variables from 1 equation.

ganeshie8 (ganeshie8):

\[3^{11 } = nk + \dfrac{k(k-1)}{2}\]

OpenStudy (vishweshshrimali5):

Okay why ?

ganeshie8 (ganeshie8):

ive just simplified..

ganeshie8 (ganeshie8):

\[2 \times 3^{11 } = 2nk + k(k-1)\]

OpenStudy (vishweshshrimali5):

Ohh okay

ganeshie8 (ganeshie8):

\[2 \times 3^{11 } = k(2n + k-1)\]

OpenStudy (vishweshshrimali5):

Okay I got that but how should I proceed after that. I am thinking of something like divisibility

OpenStudy (vishweshshrimali5):

\[\large{k|2\cdot3^{11}}\]

ganeshie8 (ganeshie8):

yes, there are (1+1)(11+1) = 24 divosors

OpenStudy (vishweshshrimali5):

Okay. But I am not going to check them one by one.. Its not fair :(

ganeshie8 (ganeshie8):

k can be any one of them, but we can reduce the number of cases significantly with few observations

OpenStudy (vishweshshrimali5):

Okay. I think we would have to use the fact that k has to be maximum.

ganeshie8 (ganeshie8):

(2n+k-1) > k

OpenStudy (vishweshshrimali5):

But that is already true...

ganeshie8 (ganeshie8):

that observation reduces the number of cases to more than half

OpenStudy (vishweshshrimali5):

Okay I would still have to check out 12 options ?

ganeshie8 (ganeshie8):

yes, this is a quadratic diaphontine equation 12 options is not too unfair :) i see we can reduce further..

OpenStudy (vishweshshrimali5):

Okay.. I too would try to get something out of it. Well thanks for reducing the options ;)

ganeshie8 (ganeshie8):

thats it, i think we're done! all 12 cases are solutions to the diaphontine equation out of these solutions, the largest possible value for k is : \(2*3^5\) since it has to be less than the other factor \(3^6\)

OpenStudy (vishweshshrimali5):

Great !! I found the solution here: http://www.artofproblemsolving.com/Wiki/index.php/1987_AIME_Problems/Problem_11 Thanks a lot @ganeshie8 .

OpenStudy (vishweshshrimali5):

I think I would have to review Diophantine Equations :)

OpenStudy (vishweshshrimali5):

Thanks again

ganeshie8 (ganeshie8):

nice :) we worked a similar problem recently @ikram002p

ganeshie8 (ganeshie8):

it is about finding the largest sequence of consecutive positive integers whose sum is 2000

OpenStudy (vishweshshrimali5):

Ohh ?

OpenStudy (vishweshshrimali5):

Well that would have lesser number of possible cases then ?

ganeshie8 (ganeshie8):

yes so much reduction is possible and can be bit tricky : \(2000 = a + (a+1) + \cdots (a+k-1)\)

ganeshie8 (ganeshie8):

but the method is exact same as earlier one... we get a diaphontine equation and solve it

OpenStudy (ikram002p):

we can also found it by finding odd factors

ganeshie8 (ganeshie8):

how does that work @ikram002p ?

OpenStudy (ikram002p):

ik how algorithm ,but dnt know the proof lol

OpenStudy (ikram002p):

i think greatest n will be ( 3^11+1)/2 then we find k from sumation